OpenStudy (anonymous):

how do you solve x^4-29x^2+100=0. Please show steps

OpenStudy (mertsj):

(x^2-25)(x^2-4)=0 (x-5)(x+5)(x-2)(x+2)=0 x=5,-5,2,-2

OpenStudy (anonymous):

thank-you!!